If two distributions have the same range, then what must be equal? a) their high and low data values 2. In a skewed distribution, the median is preferred over the mean as a measure of the center because a) the median is the most frequent number 3. If a distribution has a mean of 9 and five of the six data values are 6, 8, 11, 15, and 18, then the missing data value is: a) 0 4. Gail received a grade of 75 on her chemistry exam. Rich received a grade of 80 on his law exam. The mean for the chemistry exam is 70 with a standard deviation of 5 and the mean for the law exam is 75 with a standard deviation of 4. Whose grade is more outstanding? a) Richs, since his grade of 80 is higher than Gails 75 Questions 5-8: Given the following table of representative data values: raw score z-score percentile rank 83 95 101 1.0 5 0 1.0 2.0 2.5 40 50 75 99 53 59 65 20 (Qi) 5. Whats the missing raw score? 6. Whats the missing z-score? 7. Whats the difference between the two missing percentile ranks? 8. Since the mean is to theo 9. In a symmetric distribution, what is the approximate percentage of data values that lie between two a) 70 a) 0 of the median, this distribution isske a) left, positively standard deviations below the mean and one standard deviation above the mean? a) 84% 10. Using this set of data: (122, 134, 126, 120, 128, 130, 120, 118, 125, 122, 126, 136, 118, 122, 124, 119), what is the percentile rank of 126 rounded to the nearest percent? a) 41 11. Find the probability of getting a sum (5 or 6) when rolling a pair of dice. a) 5/18
Answer #1

1)

Range = Highest - Lowest

option D)

2)

option B) mean maybe heavily influenced by outliers

3)

missing value = 9 *6 - (6 + 8 +11 + 15 +18)

=   -4

option C) is correct

4)

z = (X - mean)/sd

z-score for Gail = (75 - 70)/5 = 1

z-score for Rich = (80 - 75)/4 = 1.25

since

Rich > Gail

option C) is correct

5)

X = mean + z* sd

83 = mean + 1 * sd

95 = mean + 2 * sd

solving mean = 71

sd = 12

z = 0

X = mean = 71

option B)

6)

z = (X - mean)/sd = (53 - 71)/12

=   -1.5

7)

percentile = 100 * P(-1.5 < Z < -0.5)

= 25 (approx)

Please rate

Please post one question at a time